# Let It Flow: Simultaneous Optimization of 3D Flow and Object Clustering

### v2

@@ Abstract
Figure 1: Performance comparison of the proposed method with self-supervised competitors on Argoverse2 Dataset [3]. Our method is able to distinguish the different motion patterns and separate objects, while other methods tend to under-segment objects and fit incorrect rigid motion. The qualitative example comes from Waymo [15] dataset.
@@ I Introduction
Accurate estimation of the 3D flow field [13, 32, 19, 12] between two consecutive point cloud scans is crucial for many applications in autonomous driving and robotics as it provides per-point motion features inherently useful for a myriad of subsequent tasks, including semantic/instance segmentation [1, 34, 14], object detection [8, 14], motion prediction [22, 4, 15] or scene reconstruction [21]. Since manually annotated data for fully-supervised training are expensive and largely not available, we focus on the self-supervised setup, which is practically useful and reaches competitive results.
@@
In the proposed self-supervised framework, we employ two losses that are visualized using the analogy of a mechanical machine [^1] Since both losses minimize the L2-norm of some quantities, we can visualize their influence on the estimated flow by ideal springs. In the ideal spring, the total conserved energy is proportional to the square of its deformation; therefore, the value of L2-loss corresponds to its energy. with springs, see Figure 2. The equilibrium of this machine corresponds to the globally optimal flow. The machine consists of two consecutive point clouds in time (black and red crosses), which are not allowed to move, flow arrows attached to blue points by swivel telescopic joints, and two losses (rigid and distance) represented by springs. The scene depicts two rigid vertical objects that are moving in different directions. We employ standard distance loss (red springs), which attracts the flow of black points toward the corresponding red points from the consecutive point cloud. Since objects in the black point cloud are close to each other, no spatial clustering would separate them correctly. In contrast to existing methods, we instead introduce overlapping clusters (blue/green/magenta regions), within which the rigidity is encouraged through rigidity loss (blue/green/magenta springs). This rigid loss constructs springs among the arrow end-points in each source rigid-clusters, which prevents the flow from deforming the rigid cluster. The color of the springs matches the color of the cluster.
@@
  • The proposed self-supervised method achieves superior accuracy among state-of-the-art methods [21, 7, 20, 17] on all tested datasets (Agroverse2 [3], Waymo [35], StereoKITTI [28])
@@ II Related Work
In the early stages of learning-based 3D scene flow estimation, methodologies relied on synthetic datasets [12, 27, 39, 26] for preliminary training. FlowNet3D [25], assimilated principles from FlowNet [13] and underwent comprehensive supervised training, employing L2 loss in conjunction with ground-truth flow annotations. BiPFN [6] orchestrated the bidirectional propagation of features from each point cloud, thereby enriching the representation of individual points. FLOT [32] introduced an innovative correspondence-based network, transforming the initial flow estimation to optimal transport estimation problem, succeeded by meticulous flow refinement through trained convolutions. The recently unveiled IHNet architecture [40] proposes an iterative hierarchical network steered by high-resolution inferred information and estimating flow in a coarse-to-fine manner.
@@
Recent methodologies have sidestepped the requirement for ground-truth flow by embracing self-supervised learning. Among the early adopters of this approach, PointPWC-Net [42] presented a wholly self-supervised methodology, combining nearest-neighbors and Laplacian losses. Similarly, in [29], a self-supervised nearest neighbor loss was implemented, ensuring the cycle consistency between forward and reverse scene flow. Najibi et al. [30] employ self-learned flow within an automated pseudo-labeling pipeline, designed for training an open-set object detector and trajectory predictor. RigidFlow [19] formulates a methodology for generating pseudo scene flow within the domain of self-supervised learning. This approach relies on piecewise rigid motion estimation applied to collections of pseudo-rigid regions, identified through the supervoxels method [24]. Flowstep3D [16] undertakes flow estimation through local and global feature matching utilizing correlation matrices and iterative warping. The method known as SCOOP [17] adopts a hybrid framework of correspondence-based feature matching coupled with a flow refinement layer employing self-supervised losses. Shen et al. [33] seamlessly integrated superpoint generation directly into their model for iterative flow refinement. Compared to supervised and self-supervised approaches, our method does not require any labeled or unlabeled training data as it is optimization-based.
Direct flow optimization
The Graph Prior approach, as introduced by [31], is rooted in pure optimization, enabling flow estimation without the need for training data. The Neural Prior[20] illustrates that flows can be optimized by incorporating a structure-based prior within the network architecture. The central objective in [21] revolves around expediting the Neural Scene Flow [21]. Li et al. identify the Chamfer distance as a computational bottleneck and use the Distance Transform as a surrogate correspondence-free loss function. The methodology outlined in Scene Flow via Distillation by [36] embodies a direct distillation framework. This innovative approach utilizes a label-free optimization method to generate pseudo-labels, which are subsequently employed to supervise a feed-forward model to achieve a better speed/performance ratio. The two newest optimization-based methods, the MBNSFP [37] and Chodosh [7] adopt Neural Prior architecture and enhanced it with spatial consistency regularization [37] and post-process cluster rigidity [7], respectively. We also use rigidity regularization, but we do not limit our method to fixed initial clusters as [37, 7]. Instead, we jointly optimize the initial cluster and flow to merge and enlarge clusters while enforcing rigidity on them progressively. Another difference between our method and the aforementioned methods is the absence of Neural Prior as structural regularization, which allows us to capture correct flows of multiple mutually-exclusive object motions.
@@
Given the presence of multiple local minima in the conventional self-supervised Chamfer distance loss, introducing regularization mechanisms becomes imperative to attain physically plausible flows—ones that faithfully trace the motion of rigid structures and objects. Implicit rigidity regularization is enforced through a robust model prior in [20, 2, 43]. The introduction of weak supervision, in the form of ego-motion and foreground segmentation, has demonstrated its capacity to provide an object-level abstraction for estimating rigid flows [11]. FlowStep3D [16] imposed constraints on the source point cloud to preserve its Laplacian when warped by the predicted flow, constituting the smoothness loss. SLIM [1] incorporated the smoothness loss into the process of learning network weights and SCOOP [17] integrated the smoothness loss within the optimization-based flow refinement layer. In [5], a novel global similarity measure takes the form of a second-order spatial compatibility measure on consensus seeds that serve as input to a weighted SVD algorithm, ultimately producing global rigid transformation. Contrary to [5], we use a spatial consistency mechanism for rigidity regularization but for multiple objects. ICP-flow [23] performs hierarchical clustering with histogram initialization for ICP registration per object to mitigate registration local minima. Such clustering usually suffers from the under-segmentation of objects. We differ from [23] by utilizing per-point soft registration and specifically addressing the under-segmentation by including the clustering in the optimization process. We also differ in performing the clustering in spatial-temporal domain, rather than geometrical space, and we do not require matching the clusters in two consecutive frames.
@@ III Method
The main goal is the self-supervised estimation of the flow field $F$ between two consecutive point clouds $P$ and $Q$. Point clouds consist of 3D points $\mathbf{p}_{i}\in P\subset\mathbb{R}^{3}$ and $\mathbf{q}_{i}\in Q\subset\mathbb{R}^{3}$. Since real-world scenes consist mostly of a static background, we follow a good practice of compensating the ego-motion [7] first before running our method. We calculate the ICP [38] as an estimate of ego motion and use it to transform the source point cloud $P$. In the rest of this section, we focus on the estimation of the remaining flow, which corresponds to the motion between transformed source point cloud $P$ and target point cloud $Q$. Consequently, the flow of point $\mathbf{p}_{i}\in P$ is a 3D vector $\mathbf{f}_{i}\in F\subset\mathbb{R}^{3}$ that is non-zero only on dynamic objects if the ego-motion is correctly estimated.
@@ III-A Method overview
The method is summarized by pseudo-code in Algorithm 1. The proposed method first initializes two sets of clusters: hard and soft. Hard rigid clusters are non-overlapping small clusters that are assumed to cover only a single rigid object or its part. Each rigid cluster $H\in\mathcal{H}$ is a small compact cluster delivered through spatial-temporal segmentation on $P$, including the adjacent temporal point clouds. In contrast, the soft rigid clusters are small overlapping clusters that are expected to overflow into neighboring objects. Each point $\mathbf{p}\in P$ is associated with the one soft cluster $S\in\mathcal{S}$, which is defined as the set of its $k$ nearest neighbors $N_{k}(\mathbf{p},P)$ from the point cloud $P$. Given these two sets of clusters, we optimize the flow to minimize (i) hard rigidity loss on hard clusters, (ii) soft rigidity loss on soft clusters, and (iii) distance loss on all points. Both rigidity losses enforce rigid flow on points within the cluster. The main difference is that the soft rigidity loss allows for outlier rejection. All losses are detailed in the following paragraphs. Once the optimized flow is available, we merge rigid clusters in $P$, whose flow goes into the same rigid cluster in $Q$. Since the algorithm is iteratively called on all consecutive pairs of point clouds, the rigidity is propagated throughout the temporal domain.
@@ III-C Hard rigidity loss
where $d_{ij}^{u}=|\mathbf{p}_{i}^{u}-\mathbf{p}_{j}^{u}|$ is the distance between the $u$-th dimension of points $\mathbf{p}_{i},\mathbf{p}_{j}\in C$ and similarly $\hat{d}_{ij}^{u}=|(\mathbf{p}_{i}^{u}+\mathbf{f}_{i}^{u})-(\mathbf{p}_{j}^{u}+\mathbf{f}_{j}^{u})|$ is the distance after applying the estimated flow on these points, and $\theta$ is a hyper-parameter of the proposed method. If the motion is rigid, the distance difference is zero, and the reward equals one; if the distance is non-zero the reward is proportionally smaller. Given this notation, we introduce hard rigidity loss
@@ III-D Soft rigidity loss
which pushes the principal eigenvalue up and consequently makes the flow on spatially compact clusters more rigid.
@@ IV Experiments
We use $k=16$ and set $\alpha$ and $\beta$ loss weights to 1. We use the learning rate of 0.004. The $\theta$ distance threshold for outlier rejection is set to 0.03 as in [5, 37]. We perform hard clustering by spatio-temporal DBSCAN over the horizon of 5-point clouds and set the epsilon parameter to 0.3 and minimal samples to 1, resulting in clustering the points to the same cluster if the Euclidian distance is lower than 0.3. The flows are optimized until a fixed number of iterations (1500) or until the convergence of the loss function is reached. For other comparative methods, we use the same optimized parameters as reported in their original papers and official implementations.
@@ IV-A Comparison with State-of-the-Art Methods
Methods Avg.$\downarrow$ Dyn.$\downarrow$ Stat.$\downarrow$ Avg.$\downarrow$ Dyn.$\downarrow$ Stat.$\downarrow$ Avg.$\downarrow$ Dyn.$\downarrow$ Stat.$\downarrow$
@@ IV-C Results on StereoKITTI benchmark
We also evaluate our method on both the full StereoKITTI and its testing subset KITTIt  following the experimental setting in [17, 29, 16]. We compare our method to the top-performing methods in self-supervised 3D scene flow and also to recent fully-supervised methods trained on FT3D [10] dataset. Our self-supervised method without any training data is on par with the most recent and top performing fully-supervised method, the IHNet [40] in terms of $AS$ (Ours 98.0% and fully-supervised IHNet 97.8%), see Figure 3.
@@
Figure 3: Comparison of the proposed method with self-supervised [17, 19, 18, 20, 31] and fully-supervised [25, 32, 40] competitors on StereoKITTI benchmark.
Figure 4: Qualitative example on Waymo Dataset. We show that by DBSCAN under-segmentation (left part of the image, where color denotes assignment to the cluster), it is impossible to separate motions of multiple objects in one cluster. When the MBNSFP [37] method converges, the flow vectors on a single cluster look uniform, resembling similar motion patterns on separate objects, while with our gradual clustering, we are able to separate motions and assign correct headings of flows. In the last row, we applied DBSCAN clustering to only flow vectors to show that we can separate the objects based on our method output flows. We could not find parameters to separate objects with MBNSFP [37] flows, as they are too close to each other.
@@ IV-D DBSCAN Clustering in Scene Flow
In order to achieve the best possible clustering in competing methods, the authors of [37, 7] tune The DBSCAN algorithm to catch the bigger objects, which is visible in parameter epsilon (distance grouping threshold) set to 0.8 and minimal number of point samples in one density-based cluster set to 30 in MBNSFP [37]. Such parametrization allows for the successful grouping of points in whole bodies of large objects such as vehicles but simultaneously groups multiple smaller objects together. Another issue stems from points that fall below such threshold and, therefore, do not form the cluster and are treated without rigidity regularization in the methods. See an example of such DBSCAN clustering in Figure 4, where the multiple pedestrians are grouped together with a tree or a vehicle in the middle of point cloud has blue points (points not assigned to any cluster, therefore without rigidity regularization).
We conclude that such clustering parametrization leads to overfitting large objects. Therefore we designed our hard clustering approach to assign all points to clusters in a "over-segmentation" manner and gradually merge the over-segmented clusters based on flow optimization and soft-clusters.

### v3

@@ Abstract
Figure 1: Performance comparison of the proposed method with self-supervised competitors on Argoverse2 Dataset [3]. Our method is able to distinguish the different motion patterns and separate objects, while other methods [7, 35, 20, 19] tend to under-segment objects and fit incorrect rigid motion. The qualitative example comes from Waymo [15] dataset.
@@ I Introduction
Accurate estimation of the 3D flow field [13, 31, 18, 12] between two consecutive point cloud scans is crucial for many applications in autonomous driving and robotics as it provides per-point motion features inherently useful for a myriad of subsequent tasks, including semantic/instance segmentation [1, 32, 14], object detection [8, 14], motion prediction [21, 4, 15] or scene reconstruction [20]. Since manually annotated data for fully-supervised training are expensive and largely not available, we focus on the self-supervised (optimization-based) setup, which is practically useful and reaches competitive results.
@@
In the proposed optimization-based framework, we employ two losses that are visualized using the analogy of a mechanical machine [^1] Since both losses minimize the L2-norm of some quantities, we can visualize their influence on the estimated flow by ideal springs. In the ideal spring, the total conserved energy is proportional to the square of its deformation; therefore, the value of L2-loss corresponds to its energy. with springs, see Figure 2. The equilibrium of this machine corresponds to the globally optimal flow. The machine consists of two consecutive point clouds in time (black and red crosses), which are not allowed to move, flow arrows attached to blue points by swivel telescopic joints, and two losses (rigid and distance) represented by springs. The scene depicts two rigid vertical objects that are moving in different directions. We employ standard distance loss (red springs), which attracts the flow of black points toward the corresponding red points from the consecutive point cloud. Since objects in the black point cloud are close to each other, no spatial clustering would separate them correctly. In contrast to existing methods, we instead introduce overlapping clusters (blue/green/magenta regions), within which the rigidity is encouraged through rigidity loss (blue/green/magenta springs). This rigid loss constructs springs among the arrow end-points in each source rigid-clusters, which prevents the flow from deforming the rigid cluster. The color of the springs matches the color of the cluster.
@@
  • The proposed optimization-based method achieves superior accuracy among state-of-the-art methods [20, 7, 19, 17] on all tested datasets (Agroverse2 [3], Waymo [33], StereoKITTI [28])
@@ II Related Work
In the early stages of learning-based 3D scene flow estimation, methodologies relied on synthetic datasets [12, 26, 37, 25, 38] for preliminary training. FlowNet3D [24], assimilated principles from FlowNet [13] and underwent comprehensive supervised training, employing L2 loss in conjunction with ground-truth flow annotations. BiPFN [6] orchestrated the bidirectional propagation of features from each point cloud, thereby enriching the representation of individual points. FLOT [31] introduced an innovative correspondence-based network, transforming the initial flow estimation to optimal transport estimation problem, succeeded by meticulous flow refinement through trained convolutions.
@@
Recent methodologies have sidestepped the requirement for ground-truth flow by embracing self-supervised learning. Among the early adopters of this approach, PointPWC-Net [41] presented a wholly self-supervised methodology, combining nearest-neighbors and Laplacian losses. Similarly, in [29], a self-supervised nearest neighbor loss was implemented, ensuring the cycle consistency between forward and reverse scene flow. RigidFlow [18] formulates a methodology for generating pseudo scene flow within the domain of self-supervised learning. This approach relies on piecewise rigid motion estimation applied to collections of pseudo-rigid regions, identified through the supervoxels method [23]. The method known as SCOOP [17] adopts a hybrid framework of correspondence-based feature matching coupled with a flow refinement layer employing self-supervised losses. Compared to supervised and self-supervised approaches, our method does not require any labeled or unlabeled training data as it is optimization-based.
Direct flow optimization and clustering
The Neural Prior[19] illustrates that flows can be optimized by incorporating a structure-based prior within the network architecture. The central objective in [20] revolves around expediting the Neural Scene Flow [20]. Li et al. identify the Chamfer distance as a computational bottleneck and use the Distance Transform as a surrogate correspondence-free loss function. The methodology outlined in Scene Flow via Distillation by [34] embodies a direct distillation framework. The two newest optimization-based methods, the MBNSFP [35] and Chodosh [7] adopt Neural Prior architecture and enhanced it with spatial consistency regularization [35] and post-process cluster rigidity [7], respectively. Aforementioned methods use DBSCAN [9] to perform clustering in single frame, whereas we use Euclidean clustering on temporally synchronized frames with bottom-up approach of connecting the points based on flows. Other related bottom-up (Agglomerative [43]) clustering approaches includes WardLinkage [39], HDBSCAN [27], BIRCH [30], which are based on mutual reachability and cluster stability. All methods use either hard clustering or hierarchical structures as a final output. They do not use overlapping clusters and flow guidance like our method.
@@
Given the presence of multiple local minima in the conventional self-supervised Chamfer distance loss, introducing regularization mechanisms becomes imperative to attain physically plausible flows—ones that faithfully trace the motion of rigid structures and objects. Implicit rigidity regularization is enforced through a robust model prior in [19, 2, 42]. The introduction of weak supervision, in the form of ego-motion and foreground segmentation, has demonstrated its capacity to provide an object-level abstraction for estimating rigid flows [11]. In [5], a novel global similarity measure takes the form of a second-order spatial compatibility measure on consensus seeds that serve as input to a weighted SVD algorithm, ultimately producing global rigid transformation. The mechanism was then extended to multi-object flow estimation in MBNSFP [35]. ICP-flow [22] performs hierarchical clustering with histogram initialization for ICP registration per object to mitigate registration local minima. Such clustering usually suffers from the under-segmentation of objects. We differ from [22] by utilizing per-point soft registration and specifically addressing the under-segmentation by including the clustering in the optimization process. We also differ in performing the clustering in spatial-temporal domain, rather than geometrical space, and we do not require matching the clusters in two consecutive frames. We also use rigidity regularization, but we do not limit our method to fixed initial clusters as [35, 7]. Instead, we jointly optimize the initial cluster (constructed with Euclidean clustering of points in spatial-temporal space) and flow to merge and enlarge clusters while enforcing rigidity on them progressively. Another difference between our method and the aforementioned methods is the absence of Neural Prior as structural regularization, which allows us to capture correct flows of multiple mutually-exclusive object motions.
@@ III Method
The main goal is the optimization-based estimation of the flow field $F$ between two consecutive point clouds $P$ and $Q$. Point clouds consist of 3D points $\mathbf{p}_{i}\in P\subset\mathbb{R}^{3}$ and $\mathbf{q}_{i}\in Q\subset\mathbb{R}^{3}$. Since real-world scenes consist mostly of a static background, we follow a good practice of compensating the ego-motion [7] first before running our method. We calculate the ICP [36] as an estimate of ego motion and use it to transform the source point cloud $P$. In the rest of this section, we focus on the estimation of the remaining flow, which corresponds to the motion between transformed source point cloud $P$ and target point cloud $Q$. Consequently, the flow of point $\mathbf{p}_{i}\in P$ is a 3D vector $\mathbf{f}_{i}\in F\subset\mathbb{R}^{3}$ describing the velocity.
@@ III-A Method overview
The method is summarized by pseudo-code in Algorithm 1. The proposed method first initializes two sets of clusters: hard and soft. Hard rigid clusters are non-overlapping small clusters that are assumed to cover only a single rigid object or its part. Each rigid cluster $H\in\mathcal{H}$ is a small compact cluster delivered through spatial-temporal segmentation on $P$, including the adjacent temporal point clouds. In contrast, the soft rigid clusters are small overlapping clusters that are expected to overflow into neighboring objects. Each point $\mathbf{p}\in P$ is associated with the one soft cluster $S\in\mathcal{S}$, which is defined as the set of its $k$ nearest neighbors $N_{k}(\mathbf{p},P)$ from the point cloud $P$. Given these two sets of clusters, we optimize the flow to minimize (i) hard rigidity loss on hard clusters, (ii) soft rigidity loss on soft clusters, and (iii) distance loss on all points. Both rigidity losses enforce rigid flow on points within the cluster. The main difference is that the soft rigidity loss allows for outlier rejection. All losses are detailed in the following paragraphs. Once the optimized flow is available, we merge rigid clusters in $P$, whose flow goes into the same rigid cluster in $Q$. Since the algorithm is iteratively called on all consecutive pairs of point clouds, the rigidity is propagated throughout the temporal domain. The flow is initialized as a zero vector, and minimization is applied as a gradient optimization to minimize the final loss function with the Adam optimizer. The reason for initialization in zero is that it is likely that the nearest neighbor is going to be the same object next time (after ICP ego-motion compensation).
@@ III-C Hard rigidity loss
where $d_{ij}^{u}=|\mathbf{p}_{i}^{u}-\mathbf{p}_{j}^{u}|$ is the distance between the $u$-th dimension of points $\mathbf{p}_{i},\mathbf{p}_{j}\in C$ and similarly $\hat{d}_{ij}^{u}=|(\mathbf{p}_{i}^{u}+\mathbf{f}_{i}^{u})-(\mathbf{p}_{j}^{u}+\mathbf{f}_{j}^{u})|$ is the distance after applying the estimated flow on these points, and $\theta$ is a hyper-parameter of the proposed method. If the motion is rigid, the distance difference is zero, and the reward equals one; if the distance is non-zero the reward is proportionally smaller. Given this notation, we introduce hard rigidity loss:
@@
Where we also clip the values of $r_{ij}$ between 0 and 1 to prevent potential numerical instability.
@@ III-D Soft rigidity loss
which pushes the principal eigenvalue up and consequently makes the flow on spatially compact clusters more rigid. The final soft rigidity is weighted by the eigenvalues of $\mathbf{A}_{m}$, that are aimed to decrease the effect of loss on outlier points.
@@ IV Experiments
We use $k=16$ and set $\alpha$ and $\beta$ loss weights to 1. We use the learning rate of 0.004 and Adam optimizer. The $\theta$ distance threshold for outlier rejection is set to 0.03 as in [5, 35]. We perform hard clustering by spatio-temporal Euclidean clustering over the horizon of 5-point clouds and group points together within the radius of 0.3m. The flows are optimized until a fixed number of iterations (1500). For other comparative methods, we use the same optimized parameters as reported in their original papers and official implementations.
@@
Figure 3: On the left - clusters from method [35], where each color denotes a single DBSCAN cluster meant for a single rigid motion fit with outlier rejection. On the right - our over-segmentation Euclidean clustering (each cluster denoted by color) and soft rigidity connections that can overflow into more objects but can be rejected as outliers. We observe that a building with three pedestrians clustered can only allow for a single rigid motion (stationary from more building points) and reject the pedestrian motion in the loss for the previous method. On the other hand, motions on the right are treated separately for each object, with the exception of two pedestrians walking right next to each other.
@@ IV-A Comparison with State-of-the-Art Methods
Methods Neural Prior Avg.$\downarrow$ Dyn.$\downarrow$ Stat.$\downarrow$ Avg.$\downarrow$ Dyn.$\downarrow$ Stat.$\downarrow$ Avg.$\downarrow$ Dyn.$\downarrow$ Stat.$\downarrow$
@@ IV-C Results on StereoKITTI benchmark
We also evaluate our method on both the full StereoKITTI and its testing subset KITTIt  following the experimental setting in [17, 29, 16]. We compare our method to the top-performing methods in self-supervised 3D scene flow and also to recent fully-supervised methods trained on FT3D [10] dataset. Our self-supervised method without any training data is on par with the most recent and top performing fully-supervised method, the IHNet [38] in terms of $AS$ (Ours 98.0% and fully-supervised IHNet 97.8%).
@@ IV-D DBSCAN Clustering in Scene Flow
In order to achieve the best possible clustering in competing methods, the authors of [35, 7] tune The DBSCAN algorithm to catch the bigger objects, which is visible in parameter epsilon (distance grouping threshold) set to 0.8 and minimal number of point samples in one density-based cluster set to 30 in MBNSFP [35]. Such parametrization allows for the successful grouping of points in whole bodies of large objects such as vehicles but simultaneously groups multiple smaller objects together. Another issue stems from points that fall below such threshold and, therefore, do not form the cluster and are treated without rigidity regularization in the methods. See an example of such DBSCAN clustering in Figure 3, where the multiple pedestrians are grouped together with a building.
Figure 4: The hard clustering parameter Epsilon which denotes the grouping radius for Euclidean clustering compared to dynamic end-point-error.
We conclude that such clustering parametrization leads to over-fitting on large objects. Therefore we designed our hard clustering approach to assign all points to clusters in a "over-segmentation" manner and gradually merge the over-segmented clusters based on flow optimization and soft-clusters, see Figure 4 for effect of grouping radius on dynamic object flow performance.